OpenStudy (anonymous):

1.) Solve the following system of equations. 3x - 2y = 6 6x - 4y = 14 (0, 0) (6, 14) Infinitely Many Solutions No Solutions 2.) What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 -7 7 9 -9 3.) Which point lies in the solution set for the following system of inequalities? y < 2x + 4 y < -2x + 2 (1, 0) (-5, -2) (0, -3) (-1, 5) Please try to help. Any help from anyone is appreciated...

OpenStudy (tennistar):

hold on , will get the answer

OpenStudy (tennistar):

-x=24 x=-24 -72-6y=12 -6y=84 y=-14 3x-2y=6 6x-4y=12 2(3x-2y=6) 6x-4y=12 6x-4y=12 infinitely many solutions

OpenStudy (tennistar):

Q1

OpenStudy (anonymous):

1. 3x_2y=6 6x-4y=14 multiply first equation with -2 0x+0y=2 then no solutions

OpenStudy (anonymous):

Uh, 2 answers for question 1... very confusing, but okay.

OpenStudy (aroub):

The third answer for question 1!

OpenStudy (aroub):

-2(3x-2y=6) 6x-4y=14 => -6x+4y=-12 6x-4y=14 ----------- 0=2 So yeah strenesmee is right :)

OpenStudy (anonymous):

Alright then...

OpenStudy (aroub):

2.) 5x + 4y = 1 4x + 3y = -1 -3(5x+4y=1) 4(4x+3y=-1) => -15x-12y=-3 16x+12y=-4 -------------- x=-7 substitute x in any of the equations to find y: 4(-7)+3y=-1 -28+3y=-1 3y=27 y=9

OpenStudy (aroub):

Any doubts?

OpenStudy (anonymous):

Nope, none at all ^_^

OpenStudy (aroub):

=D

OpenStudy (anonymous):

Would you happen to know anything about the last question on here? ----^

OpenStudy (aroub):

Oh, the third question is graphing x_x

OpenStudy (anonymous):

If you can;t get it, it's fine. I'm thankful for all the help you've already given me :)

OpenStudy (aroub):

Umm, no it's okay. I'm going to try to graph it here.

OpenStudy (aroub):

 3.) y < 2x + 4 y < -2x + 2 For every inequality there is an associated equation for it. y < 2x + 4 = y=2x+4 => (0,4) and (-2,0) y < -2x + 2 = y=-2x+2 => (0,2) and (1,0)